OpenStudy (anonymous):

A Social Security number has nine digits, using the digits 0 through 9 with repetition of digits allowed. How many Social Security numbers have at least one repeated digit?

OpenStudy (anonymous):

We can approach this from the other end. If the SS number were ten digits, then we would know that the numbers 0-9 could be used once with 10! permutations. We would subtract this from 10^10 possible ten-digit numbers to get those that had one or more repetition. If we had only 1-9 to use, we could find the SS repeats similarly from 10^9 - 9!. My problem is we have 10 digits for the nine-digit number. Let's use all but 0, then all but 1, then all but 2, etc. We would have 9! possibilities times each one of these, so 10*9! = 10!. Perhaps. My best estimate: 10^9 - 10! = # SS numbers with one or more repetitions.

Parth (parthkohli):

Find the number of Social Security numbers that have no repeated digits.

OpenStudy (anonymous):

Yes, that's what I attempted. Then the answer to the question would be all possible minus this.

OpenStudy (anonymous):

Welcome to OpenStudy @nm61101 ! Here's a link to help you around OpenStudy! http://openstudy.com/study#/updates/52f5a913e4b0512c88e5ef0f

OpenStudy (anonymous):

no digits repeated is (10P9) so 10^9 - (10P9) = 996371200

OpenStudy (anonymous):

@sourwing Nice to have my guess confirmed. Thanks.

OpenStudy (anonymous):

Thank you!